# Improving Autoformalization Using Direct Dependency Retrieval

### v1

@@ 3.2 Suffix Array Auxiliary
To accurately and efficiently verify the existence of dependency objects generated by LLMs within a formal library, we introduce a checking method based on suffix arrays (liu2024infini). The core idea of this approach is to first concatenate the identifiers of all defined objects in the library, such as pre-proven theorems, into a single long string using a specific delimiter. Subsequently, a suffix array is constructed from this concatenated long string in linear complexity (lee2022deduplicating; karkkainen2006linear). By definition, a suffix array is an array of integers that stores the starting indices of all suffixes of the original string, sorted in lexicographical order (karkkainen2006linear). Figure 2 illustrates the suffix array auxiliary check process with a simplified example, assuming the library contains only two objects and there are three pending items. The sorted nature of the suffix array facilitates efficient matching of these pending items via binary search. Specifically, for each pending item, we can locate the range of all suffixes that have the item as a prefix within the suffix array. This process enables the rapid retrieval of all library objects corresponding to the pending items, thereby completing their existence verification.
@@
The suffix array auxiliary check is a highly efficient method for matching and verification, capable of meeting the real-time requirements of large-scale requests (liu2024infini). For a formal library containing $N$ objects with an average length of $d$, and $M$ pending items with an average length of $s$, conventional string matching methods typically involve brute-force comparison, resulting in a time complexity of $\mathcal{O}\left((s+d)MN\right)$. In contrast, the suffix array auxiliary check operates on a suffix array constructed over the entire library. For each pending item, it utilizes binary search on the suffix array to locate matches. The overall time complexity of this approach is $\mathcal{O}(sM\log N)$. This significant reduction in computational complexity enables real-time responsiveness, thereby making the method suitable for deployment in online application systems (liu2024infini).
@@ 3.3 Constructing High-quality Training Dataset
Fine-tuning an accurate and efficient DDR model necessitates a dataset composed of pairs of mathematical statements and their corresponding library dependencies. However, no such public dataset is readily available for this purpose. The primary challenge lies in the difficulty of accurately extracting and efficiently verifying library dependencies from existing formal statements. To address this, we utilize the suffix array auxiliary check to process large-scale autoformalizer datasets and construct the required training data. We chose FineLeanCorpus (peng2025criticlean) as our source, a dataset containing over 500,000 samples that span a wide range of mathematical domains and difficulty levels. After partitioning the data into training and test sets based on problem difficulty, we consolidated all training data and filtered the potential library dependencies. Leveraging the high efficiency of our suffix array-based approach, we processed the entire dataset in under two hours, resulting in a dataset suitable for fine-tuning the DDR model. This resulting dataset has been made open-source at https://huggingface.co/datasets/Palca/FineLeanCorpus_DDR. We anticipate that this contribution will foster further research on RAG-based approaches within the field of formal mathematical proving.
Table 2: Comparison of our DDR with baseline methods. R@k denotes the topological-informalization retrieval method from liu2025rethinking, where the top-k retrieved results are used as dependency outputs. The symbol $\uparrow$ indicates that our DDR consistently outperforms the baselines.
Difficulty R@5 R@10 DDR Precision Recall F1 $Precision\uparrow Recall\uparrow F1\uparrow$ Diff01 $0.033$ 0.093 333 30.093\,333\,3 0.046 424 963 924 963 930.046\,424\,963\,924\,963\,93 $0.023$ 0.132 916 666 666 666 660.132\,916\,666\,666\,666\,66 0.037 839 452 214 452 2160.037\,839\,452\,214\,452\,216 $0.845$ 0.737 250.737\,25 0.766 428 571 428 571 30.766\,428\,571\,428\,571\,3 Diff23 $0.083$ 0.174 601 190 476 190 470.174\,601\,190\,476\,190\,47 0.104 828 255 078 255 080.104\,828\,255\,078\,255\,08 $0.053$ 0.233 892 857 142 857 120.233\,892\,857\,142\,857\,12 0.081 945 344 687 991 760.081\,945\,344\,687\,991\,76 $0.925$ 0.580 406 746 031 7460.580\,406\,746\,031\,746 0.664 218 253 968 2540.664\,218\,253\,968\,254 Diff45 $0.057$ 0.131 416 666 666 666 650.131\,416\,666\,666\,666\,65 0.074 401 875 901 875 910.074\,401\,875\,901\,875\,91 $0.0375$ 0.173 416 666 666 666 660.173\,416\,666\,666\,666\,66 0.058 731 601 731 601 740.058\,731\,601\,731\,601\,74 $0.91$ 0.683 249 999 999 999 90.683\,249\,999\,999\,999\,9 0.736 361 471 861 471 80.736\,361\,471\,861\,471\,8 Diff67 $0.087$ 0.186 214 285 714 285 750.186\,214\,285\,714\,285\,75 0.110 370 129 870 129 880.110\,370\,129\,870\,129\,88 $0.068$ 0.296 952 380 952 380 90.296\,952\,380\,952\,380\,9 0.104 312 102 113 572 710.104\,312\,102\,113\,572\,71 $0.92$ 0.569 555 555 555 555 60.569\,555\,555\,555\,555\,6 0.651 988 095 238 095 30.651\,988\,095\,238\,095\,3 Diff89 $0.115$ 0.193 134 920 634 920 630.193\,134\,920\,634\,920\,63 0.134 166 527 916 527 940.134\,166\,527\,916\,527\,94 $0.0755$ 0.257 095 238 095 238 10.257\,095\,238\,095\,238\,1 0.110 275 899 109 490 430.110\,275\,899\,109\,490\,43 $0.855$ 0.462 984 126 984 1270.462\,984\,126\,984\,127 0.549 760 442 773 600 70.549\,760\,442\,773\,600\,7
@@ 4 Direct Dependency Retrieval-augmented Autoformalization
While the current autoformalization paradigm approach effectively prevents the hallucination of formal objects that do not exist within the library (liu2025rethinking; yang2023leandojo; azerbayev2023proofnet), the accuracy of existing retrieval methods is limited (liu2025rethinking). Furthermore, when the number of retrieved items exceeds the actual number of required dependencies, these superfluous items introduce noise into the autoformalization process. Consequently, such methods typically require prior knowledge of the number of dependencies to mitigate this noise, a prerequisite that is often impractical in application. DDR achieves significantly higher precision and recall in identifying the actual dependencies from Mathlib. It operates without requiring prior knowledge of the number of dependencies and better approximates the distribution of true dependencies, leading to superior generalization capabilities. Comprehensive experimental results validate the effectiveness and superiority of the DDR method.
Table 3: Comparison and ablation results of retrieval-augmented methods. Above the dashed line are the ICL methods; below are the SFT methods. All ICL methods use the same prompt as in liu2025rethinking. MMA denotes the LLM fine-tuned on MMA’s Lean subset. RA denotes the pure autoformalizer method in liu2025rethinking. R@5 and R@10 denote retrieval-augmented methods using top-5 and top-10 retrieved results, respectively. RA+DDR is our main method.
@@
Method    Diff01    Diff23    Diff45    Diff67    Diff89 TypeC BEq TypeC BEq TypeC BEq TypeC BEq TypeC BEq DeepSeek-R1 0.502 512 562 814 070 30.502\,512\,562\,814\,070\,3 0.216 080 402 010 050 240.216\,080\,402\,010\,050\,24 $0.255$ $0.17$ 0.277 777 777 777 777 80.277\,777\,777\,777\,777\,8 0.176 767 676 767 676 770.176\,767\,676\,767\,676\,77 0.272 727 272 727 272 70.272\,727\,272\,727\,272\,7 0.191 919 191 919 191 90.191\,919\,191\,919\,191\,9 0.217 171 717 171 717 180.217\,171\,717\,171\,717\,18 0.166 666 666 666 666 660.166\,666\,666\,666\,666\,66 Ling-flash-2.0 $0.635$ $0.575$ $0.365$ $0.33$ $0.425$ $0.41$ $0.305$ $0.3$ $0.17$ $0.16$ Claude-3.5-Sonnet $0.67$ $0.61$ $0.52$ $0.48$ $0.55$ $0.53$ $0.405$ $0.395$ $0.335$ $0.315$ GPT-4o $0.66$ $0.56$ $0.49$ $0.48$ $0.52$ $0.515$ $0.44$ $0.435$ $0.265$ $0.265$ Qwen2.5-Max $0.645$ $0.565$ $0.4$ $0.385$ $0.505$ $0.505$ $0.4$ $0.4$ $0.25$ $0.235$ MMA $0.16$ $0.145$ $0.185$ $0.17$ $0.305$ $0.295$ $0.21$ $0.21$ $0.15$ $0.15$ RA $0.245$ $0.21$ 0.441 340 782 122 9050.441\,340\,782\,122\,905 0.413 407 821 229 050 30.413\,407\,821\,229\,050\,3 0.497 206 703 910 614 50.497\,206\,703\,910\,614\,5 0.491 620 111 731 843 560.491\,620\,111\,731\,843\,56 0.388 888 888 888 888 90.388\,888\,888\,888\,888\,9 0.382 716 049 382 716 030.382\,716\,049\,382\,716\,03 $0.285$ $0.275$ RA+R@5 $0.325$ $0.29$ $0.535$ $0.5$ $0.535$ $0.53$ $0.465$ $0.455$ $0.36$ $0.33$ RA+R@10 $0.335$ $0.3$ $0.48$ $0.45$ $0.545$ $0.535$ $0.475$ $0.455$ $0.31$ $0.31$ RA+DDR $0.36$ $0.315$ $0.56$ $0.52$ $0.60$ $0.59$ $0.54$ $0.51$ $0.41$ $0.38$
4.1 Dataset Overview
To construct and evaluate our proposed DDR method, we used the FineLeanCorpus (peng2025criticlean), a dataset that comprises a vast collection of natural language mathematical statements paired with their corresponding formalizations in Lean 4. To prevent data leakage from potential overlap with other Lean 4 datasets (due to its diverse origins), we split the corpus into training and test sets based on its intrinsic difficulty attribute.
@@ 4.1 Dataset Overview
The difficulty attribute is a key metric in FineLeanCorpus designed to quantify the challenge of formalizing informal natural language mathematical statements into Lean 4 statements, rated on an 11-point scale from 0 to 10. In our experimental setup, we consolidated samples with a difficulty of 10 into the difficulty 9 category.
A notable observation is that samples categorized as difficulty 0 and difficulty 1 present a counter-intuitively high formalization challenge. This is primarily because the natural language statements in this category often do not represent abstract descriptions of the problem’s mathematical essence, as illustrated by examples in appendix A. Our subsequent experimental results corroborate this finding: for tasks of difficulty 0 and difficulty 1, the performance of ICL significantly surpasses that of SFT models, a phenomenon that aligns with the analysis.
We applied the suffix algorithm described in Section 3.2 to extract Mathlib dependencies for all statements within each difficulty category. The statistical results are summarized in Table 1. The data reveals a positive correlation between a statement’s general difficulty and its degree of Mathlib dependency: categories of higher difficulty typically exhibit both a higher dependency rate and a greater number of dependencies. This finding aligns with our intuition and preliminary observations of the dataset.
@@
For the experimental setup, we constructed five test sets. Specifically, we randomly sampled 100 statements from each difficulty category and then combined the samples from every two adjacent difficulty categories. This process resulted in five test sets, named Diff01, Diff23, Diff45, Diff67, and Diff89, where Diffij denotes the combined set of samples from difficulty levels i and j, with each test set containing 200 statements. All remaining samples in the dataset were used to fine-tune the DDR model.
@@ 4.2 Setup
We employ precision, recall, and F1 score as quantitative metrics to evaluate the performance of our dependency retrieval results. A key challenge in this evaluation arises from the discrepancy between the outputs of DDR, which directly generate dependency names, and the ground-truth labels, which are Mathlib object identifiers extracted from source code. This leads to name ambiguity, as source code often contains unqualified or partially qualified names. To accurately identify true tositives for our evaluation, we introduce a suffix-based matching criterion. Specifically, we represent each identifier as a sequence of dot-separated components. A retrieved identifier is considered to match a ground-truth label if the component sequence of one is a suffix of the other.
To evaluate the semantic equivalence between the formal statements predicted by the model and the ground-truth formal statements, we employ two criteria: TypeC (Type Checking) and BEq (Bidirectional Extended Definitional Equality) (liu2025rethinking). TypeC ensures the syntactic correctness of expressions, whereas BEq is utilized to determine their semantic equivalence. Building upon these criteria, we define the TypeC@k and BEq@k evaluation metrics. These metrics measure the percentage of problems for which at least one of the top-k predicted statements is deemed equivalent to the ground-truth statement under the corresponding criterion.
Figure 3: Comparison of our DDR with baselines. (a)–(c) show results for R@5, R@10, and DDR, respectively. R@k is the topological-informalization retrieval method from liu2025rethinking. Light and wider bars indicate the mean; dark and thinner bars show the standard deviation. Gray: F1; Red: Recall; Green: Precision.
To comprehensively evaluate our proposed method, we established several baseline models and conducted ablation studies. For the ablation studies, we employed the RA (RAutoformalizer) from liu2025rethinking as a pure formalizer baseline, and RA+R@k as an autoformalizer that incorporates the top-k retrieved dependencies using the method in liu2025rethinking. The RA was fine-tuned based on InternLM2-Math-Base-7B (ying2024internlm). Our baselines cover multiple aspects: first, we evaluated a range of state-of-the-art general LLMs via ICL, including DeepSeek-R1 (guo2025deepseek), Ling-flash-2.0 (ai2025ming), Claude-3.5-Sonnet (anthropic_claude35_sonnet_2024), GPT-4o (achiam2023gpt), and Qwen2.5-Max (qwen2025qwen25technicalreport); second, we included an InternLM2-Math-Base-7B (ying2024internlm) model fine-tuned on the MMA Lean subset (jiang2023multilingual). For our proposed DDR method, we selected Qwen3-32B (yang2025qwen3) as its base model. This choice highlights a core advantage of DDR: its flexibility in selecting a base dependency retrieval model without being restricted to specific embedding models.
During the decoding stage for all experiments, the temperature was set to 0.7, and we generated 8 samples for each problem. The first generated output was used to calculate Metric@1, while all 8 outputs were used for Metric@8. To ensure a fair comparison, the dependency retrieval step for the DDR method was performed only once, and its results were reused for all subsequent Lean 4 statement formalization generations. It is noteworthy that none of the autoformalization models discussed in the main text, including the fine-tuned ones, have undergone SFT on the FineLeanCorpus (peng2025criticlean). Further details on the experimental setup are provided in appendix B.

### v2

@@ 3.2 Suffix Array Check
To accurately and efficiently verify the existence of dependency objects generated by LLMs within a formal library, we introduce a checking method based on suffix arrays (liu2024infini). The core idea of this approach is to first concatenate the identifiers of all defined objects in the library, such as pre-proven theorems, into a single long string using a specific delimiter. Subsequently, a suffix array is constructed from this concatenated long string in linear complexity (lee2022deduplicating; karkkainen2006linear). By definition, a suffix array is an array of integers that stores the starting indices of all suffixes of the original string, sorted in lexicographical order (karkkainen2006linear). Figure 2 illustrates the SAC process with a simplified example, assuming the library contains only two objects and there are three pending items. The sorted nature of the suffix array facilitates efficient matching of these pending items via binary search. Specifically, for each pending item, we can locate the range of all suffixes that have the item as a prefix within the suffix array. This process enables the rapid retrieval of all library objects corresponding to the pending items, thereby completing their existence verification.
@@
SAC is an efficient method for verification, capable of meeting the real-time requirements of large-scale requests (liu2024infini). For a formal library containing $N$ objects with an average length of $d$, and $M$ pending items with an average length of $s$, conventional string matching methods typically involve brute-force comparison, resulting in a time complexity of $\mathcal{O}\left((s+d)MN\right)$. In contrast, the suffix array auxiliary check operates on a suffix array constructed over the entire library. For each pending item, it utilizes binary search on the suffix array to locate matches. The overall time complexity of this approach is $\mathcal{O}(sM\log N)$. This significant reduction in computational complexity enables real-time responsiveness (liu2024infini).
@@ 3.3 Constructing High-quality Training Dataset
Fine-tuning a precise and efficient DDR model necessitates a dataset of pairs, each comprising a mathematical statement and its corresponding library dependencies. However, no such public dataset is currently available, and its construction is primarily challenged by the difficulty of accurately extracting and efficiently verifying library dependencies from existing formal statements. To address this, we employ the above SAC method to process a large-scale autoformalization corpus to generate the required training data. We select FineLeanCorpus (peng2025criticlean) as the source corpus, which contains over 500,000 samples spanning a wide range of mathematical domains and difficulty levels. After partitioning the corpus into training and test sets based on problem difficulty, we leverage the efficiency of SAC to process the entire training partition in under two hours. This process yield a high-quality dataset suitable for fine-tuning the DDR model. This resulting dataset has been made open-access at https://huggingface.co/datasets/Palca/FineLeanCorpus_DDR. We anticipate that this contribution will foster further research on the field of formal mathematical proving.
@@ 4 Direct Dependency Retrieval-augmented Autoformalization
Method    Diff01    Diff23    Diff45    Diff67    Diff89 Precision Recall Precision Recall Precision Recall Precision Recall Precision Recall Claude-3.5-Sonnet 0.490 249 999 999 999 960.490\,249\,999\,999\,999\,96 $0.5235$ 0.277 160 714 285 714 260.277\,160\,714\,285\,714\,26 0.384 761 904 761 904 750.384\,761\,904\,761\,904\,75 $0.214$ 0.340 333 333 333 333 30.340\,333\,333\,333\,333\,3 0.279 833 333 333 333 30.279\,833\,333\,333\,333\,3 0.434 710 317 460 317 40.434\,710\,317\,460\,317\,4 0.272 853 354 978 354 950.272\,853\,354\,978\,354\,95 0.333 990 079 365 079 40.333\,990\,079\,365\,079\,4 DeepSeek-R1 0.531 833 333 333 333 30.531\,833\,333\,333\,333\,3 0.557 833 333 333 333 30.557\,833\,333\,333\,333\,3 0.240 378 663 003 6630.240\,378\,663\,003\,663 0.353 684 523 809 523 750.353\,684\,523\,809\,523\,75 0.193 174 603 174 603 170.193\,174\,603\,174\,603\,17 0.323 297 619 047 6190.323\,297\,619\,047\,619 0.230 887 371 615 312 780.230\,887\,371\,615\,312\,78 0.368 638 888 888 888 850.368\,638\,888\,888\,888\,85 0.262 246 031 746 031 70.262\,246\,031\,746\,031\,7 0.371 527 777 777 777 80.371\,527\,777\,777\,777\,8 GPT-4o 0.538 750 000 000 000 10.538\,750\,000\,000\,000\,1 0.548 083 333 333 333 40.548\,083\,333\,333\,333\,4 0.282 880 952 380 952 40.282\,880\,952\,380\,952\,4 0.317 470 238 095 238 10.317\,470\,238\,095\,238\,1 0.203 500 000 000 000 010.203\,500\,000\,000\,000\,01 0.247 499 999 999 999 970.247\,499\,999\,999\,999\,97 0.262 416 666 666 666 70.262\,416\,666\,666\,666\,7 0.318 698 412 698 412 70.318\,698\,412\,698\,412\,7 0.274 305 555 555 555 50.274\,305\,555\,555\,555\,5 0.289 980 158 730 158 750.289\,980\,158\,730\,158\,75 Ling-flash-2.0 0.449 250.449\,25 0.469 333 333 333 333 40.469\,333\,333\,333\,333\,4 0.173 851 190 476 190 50.173\,851\,190\,476\,190\,5 0.236 178 571 428 571 430.236\,178\,571\,428\,571\,43 0.147 541 666 666 666 650.147\,541\,666\,666\,666\,65 0.219 499 999 999 999 950.219\,499\,999\,999\,999\,95 0.214 780 122 655 122 670.214\,780\,122\,655\,122\,67 0.319 948 412 698 412 70.319\,948\,412\,698\,412\,7 0.181 663 295 038 2950.181\,663\,295\,038\,295 0.247 426 587 301 587 310.247\,426\,587\,301\,587\,31 Qwen-Max 0.479 464 285 714 285 70.479\,464\,285\,714\,285\,7 0.511 250.511\,25 0.263 158 730 158 730 150.263\,158\,730\,158\,730\,15 0.293 380 952 380 952 40.293\,380\,952\,380\,952\,4 0.227 833 333 333 333 30.227\,833\,333\,333\,333\,3 0.276 583 333 333 333 350.276\,583\,333\,333\,333\,35 0.262 250.262\,25 0.331 380 952 380 952 40.331\,380\,952\,380\,952\,4 0.266 392 857 142 857 10.266\,392\,857\,142\,857\,1 0.323 132 936 507 936 450.323\,132\,936\,507\,936\,45 R@5 0.022 000 000 000 000 0020.022\,000\,000\,000\,000\,002 0.069 166 666 666 666 650.069\,166\,666\,666\,666\,65 $0.073$ 0.150 970 238 095 238 070.150\,970\,238\,095\,238\,07 0.048 999 999 999 999 9950.048\,999\,999\,999\,999\,995 0.119 250 000 000 000 010.119\,250\,000\,000\,000\,01 $0.079$ 0.171 166 666 666 666 660.171\,166\,666\,666\,666\,66 $0.085$ 0.143 722 222 222 222 220.143\,722\,222\,222\,222\,22 R@10 0.016 500 000 000 000 0040.016\,500\,000\,000\,000\,004 0.100 416 666 666 666 650.100\,416\,666\,666\,666\,65 0.044 499 999 999 999 990.044\,499\,999\,999\,999\,99 0.192 136 904 761 904 760.192\,136\,904\,761\,904\,76 $0.032$ 0.150 583 333 333 333 320.150\,583\,333\,333\,333\,32 0.059 500 000 000 000 0040.059\,500\,000\,000\,000\,004 0.265 476 190 476 190 530.265\,476\,190\,476\,190\,53 $0.061$ 0.206 480 158 730 158 70.206\,480\,158\,730\,158\,7 DDR 0.85 0.74 0.92 0.58 0.91 0.68 0.91 0.57 0.82 0.44
Table 3: Comparison of the retrieval dependency results. Results shown are filtered (i.e., without hallucinations).
Figure 3: LLM consistency across multiple attempts. Subplots (a–e) correspond to the performance of DeepSeek-R1 on datasets Diff01, Diff23, Diff45, Diff67, and Diff89, respectively. Each subplot illustrates the proportion of problems ($y$ axis) for which at least k (from 1 to 8, $x$ axis) out of 8 attempts successfully pass the BEq verification. Blue: DDR; Orange: ICL; Green: N/A.
This section presents a comprehensive evaluation of the proposed method. Our evaluation is two-fold: first, we directly assess the performance of DDR on the dependency retrieval task. Second, we measure the improvements it brings by evaluating the pass@8 success rate on the downstream autoformalization task. Furthermore, the experiments provide additional insights, such as a discussion on hallucinations in ICL methods and an analysis of the difficulty of autoformalization datasets.
4.1 Dataset Overview and Difficulty Analysis
@@
The difficulty attribute is a key metric in FineLeanCorpus designed to quantify the challenge of formalization, rated on an 11-point scale from 0 to 10. In our experimental setup, we consolidated samples with a difficulty of 10 into the difficulty 9 category. The statistical results of Mathlib dependency on datasets are summarized in Table 1.
@@
We randomly sampled 100 statements from each difficulty category and then combined the samples from every two adjacent difficulty categories as the test set. This process resulted in five test sets, named Diff01, Diff23, Diff45, Diff67, and Diff89, where Diffij denotes the combined set of samples from difficulty levels i and j, with each test set containing 200 statements. All remaining samples in the dataset were used to fine-tune the DDR model.
The original difficulty ratings for the FineLeanCorpus (peng2025criticlean) datasets, assigned by an LLM grader, exhibit a notable discrepancy with our experimental results in Table 4 and Table 5. To investigate this inconsistency, we propose assessing the actual formalization difficulty using two alternative metrics: Depend Rate and Depend Length. Based on the data in Table 1, we can calculate the weighted averages of Depend Rate and Depend Length for the Diff23, Diff45, Diff67, and Diff89 subsets. The results are as follows: Diff45 has the lowest values (0.662, 1.545), the values for Diff23 (0.792, 1.937) and Diff67 (0.786, 2.197) are higher, and Diff89 has the highest values (0.822, 2.445). These metrics reveal a difficulty ranking of Diff89 > Diff67 > Diff23 > Diff45. This ranking aligns closely with our experimental observations, suggesting that Depend Rate and Depend Length, derived via our DDR and SAC, can serve as a potential indicator of formalization difficulty along with LLM-based graders. The Diff01 subset is excluded from this analysis because its imprecise informal descriptions introduce extraneous formalization challenges, making its difficulty hard to assess with these metrics. Detailed discussion are presented in appendix B
Figure 4: LLM efficiency across multiple attempts. Subplots (a–e) correspond to the performance of DeepSeek-R1 on datasets Diff01, Diff23, Diff45, Diff67, and Diff89, respectively. Each subplot illustrates the intermediate pass@k score ($y$ axis) out of 8 attempts using BEq verification, k range from 1 to 8 ($x$ axis). Blue: DDR; Orange: ICL; Green: N/A.
Method    Diff01    Diff23    Diff45    Diff67    Diff89 TypeC BEq TypeC BEq TypeC BEq TypeC BEq TypeC BEq MMA $0.47$ $0.03$ $0.455$ $0.075$ $0.6$ $0.125$ $0.5$ $0.045$ $0.385$ $0.03$ RA $0.67$ $0.04$ $0.74$ $0.13$ $0.77$ $0.22$ $0.73$ $0.085$ $0.61$ $0.03$ RA+R@5 0.880 0.070 0.900 $0.125$ $0.895$ $0.21$ $0.855$ $0.095$ $0.765$ $0.05$ RA+R@10 $0.815$ $0.045$ $0.86$ $0.115$ $0.84$ $0.215$ $0.825$ $0.085$ $0.73$ $0.055$ RA+DDR $0.87$ 0.070 0.900 0.140 0.930 0.255 0.865 0.105 0.810 0.065
Table 4: Comparison of the autoformalization results using SFT methods. MMA denotes the LLM fine-tuned on MMA’s Lean subset. RA refers to the pure autoformalization method from liu2025rethinking, while R@5 and R@10 denote its retrieval-augmented variants using the top-5 and top-10 retrieved results, respectively. RA+DDR denotes the new retrieval-augmented variant incorporating the proposed DDR.
@@ 4.2 Setup
Due to space constraints, we defer the detailed description of our experimental setup to the appendix C. Our evaluation utilizes two primary metrics: Type Checking (TypeC) and Bidirectional Extended Definitional Equality (BEq). We compare our method against a comprehensive set of baseline methods, including MMA (jiang2023multilingual), DeepSeek-R1 (guo2025deepseek), Ling-flash-2.0 (ai2025ming), Claude-3.5-Sonnet (anthropic_claude35_sonnet_2024), GPT-4o (achiam2023gpt), Qwen-Max (qwen2025qwen25technicalreport), as well as RA+R@5 and RA+R@10 (liu2025rethinking).
4.3 Hallucination Study
Prior to the introduction of DDR, a sound mechanism for assessing dependency hallucinations in the outputs of LLMs on autoformalization tasks was lacking. The proposed DDR framework, through the integrated SAC, provides a practical and efficient solution to this problem. This method determines whether each generated dependency exists within the formal library, thereby enabling the quantification of hallucination severity. Define the hallucination rate (Hall) as the proportion of hallucinatory items in a given set of dependencies. For instance, if a model outputs (Real.div, Real.max, Nat.succ) and only Nat.succ is authentic, the Hall is 2/3. $\textbf{Hall}_{\textbf{m}}$ and $\textbf{Hall}_{\textbf{std}}$ denote the mean and standard deviation of this rate across the test set, respectively. The hallucination rates for various mainstream dependency retrieval methods are presented in Table 2.
The analysis reveals that R@5 and R@10, as select-based methods, are inherently free from hallucination by design, since their outputs are drawn directly from the formal library. In contrast, all generative retrieval methods are susceptible to hallucination. Our proposed DDR method exhibits outstanding performance, with a mean hallucination rate ($\textbf{Hall}_{\textbf{m}}$) below 0.02 and a standard deviation ($\textbf{Hall}_{\textbf{std}}$) below 0.09, indicating nearly non-existent hallucination. In stark contrast, other ICL-based methods suffer from severe hallucination, with both the mean and standard deviation of their Hall values hovering around 0.30. This implies that, on average, approximately 30% of the dependencies generated by these methods are hallucinatory.
Method RAG    Diff01    Diff23    Diff45    Diff67    Diff89 TypeC BEq TypeC BEq TypeC BEq TypeC BEq TypeC BEq Claude-3.5-Sonnet N/A $0.845$ $0.14$ $0.755$ $0.21$ 0.800 $0.275$ 0.680 0.130 $0.53$ $0.035$ Claude-3.5-Sonnet ICL $0.74$ $0.125$ $0.515$ $0.15$ $0.65$ $0.24$ $0.46$ $0.075$ $0.385$ $0.04$ Claude-3.5-Sonnet DDR 0.905 0.165 0.790 0.235 $0.77$ 0.305 $0.63$ 0.130 0.635 0.055 DeepSeek-R1 N/A $0.91$ $0.1$ $0.745$ $0.125$ $0.82$ $0.27$ $0.78$ $0.115$ $0.66$ $0.085$ DeepSeek-R1 ICL $0.885$ $0.11$ $0.76$ $0.22$ $0.845$ $0.34$ $0.76$ $0.145$ $0.575$ $0.08$ DeepSeek-R1 DDR 0.940 0.170 0.875 0.280 0.950 0.395 0.875 0.180 0.705 0.090 GPT-4o N/A $0.905$ $0.1$ $0.765$ 0.180 $0.8$ $0.175$ $0.75$ $0.09$ $0.575$ $0.05$ GPT-4o ICL $0.88$ $0.1$ $0.675$ $0.14$ $0.73$ $0.185$ $0.63$ $0.06$ $0.405$ $0.05$ GPT-4o DDR 0.940 0.125 0.775 0.180 0.850 0.280 0.780 0.100 0.630 0.060 Ling-flash-2.0 N/A 0.855 $0.105$ 0.595 0.150 0.670 $0.235$ $0.51$ $0.09$ $0.38$ 0.055 Ling-flash-2.0 ICL $0.775$ $0.075$ $0.42$ $0.1$ $0.6$ $0.225$ $0.395$ $0.08$ $0.295$ 0.055 Ling-flash-2.0 DDR $0.775$ 0.115 0.595 0.150 $0.64$ 0.260 0.535 0.105 0.440 0.055 Qwen-Max N/A $0.795$ $0.095$ $0.605$ 0.155 0.675 0.270 $0.62$ $0.11$ $0.44$ $0.060$ Qwen-Max ICL $0.73$ $0.065$ $0.445$ $0.105$ $0.555$ $0.17$ $0.445$ $0.075$ $0.35$ $0.060$ Qwen-Max DDR 0.835 0.110 0.665 $0.15$ $0.575$ $0.21$ 0.695 0.115 0.545 0.065
Table 5: Comparison of the autoformalization results using ICL methods. RAG column indicates the dependency retrieval strategy used in autoformalization: N/A denotes no retrieval, ICL denotes prompt-based retrieval, and DDR refers to the proposed method.